Hello,
I'm to find the third roots of 343i written in polar form with arguments in radians. I don't know how to approach this one. Is it actually "0 + 343i"? If so, I find the absolute value to be 343. Then tan (b/a) = tan (343/0) which is undefined. How do I find the angles?
Please help me as I'm lost...

Answer by Alan3354(6088)

I'm to find the third roots of 343i written in polar form with arguments in radians. I don't know how to approach this one. Is it actually "0 + 343i"? If so, I find the absolute value to be 343. Then tan (b/a) = tan (343/0) which is undefined. How do I find the angles?
--------------
It is 0 + 343i, and r = 343.
The angle is 90, but you don't need the tangent
z = 0 + 343i = 343cis90
--------
Cube roots are:
7cis30 = 7cis(pi/6) = 7cos(30) + 7isin(30) = 7sqrt(3)/3 + 3i/2
---------------------
Do these 2 the same way.
7cis150 = 7cis(5pi/6)
7cis270 = 7cis(2pi/3)